When the roller derby car reaches the bottom of the hill. How fast is it traveling (point C)? (using g = 10m/s/s and create an LOL to help solve the problem)
1
Expert's answer
2020-05-18T10:16:24-0400


From the LOL diagram above, we see that the potential energy turned into work done against friction and kinetic energy of derby car motion:


"PE=FR+KE,\\
mgh=\mu mgx+\frac{mv^2}{2},\\
\space\\
v=\sqrt{2g(h-\mu x)}."

Just substitute the initial height "h" , coefficient of friction "\mu", and the distance "x" where the friction was involved.